OpenStudy (anonymous):

If a sample of pure calcium (Ca) contains 4.69 x 1022 atoms of iodine, how many moles of calcium are in the sample? 0.0779 mol 0.130 mol 7.78 mol 12.8 mol

OpenStudy (anonymous):

if you look at the iodine as impurity than you cant tell what amount of Ca you have cause there cant be established any relationship between impurity and Ca, only if you had abundance of impurity in Ca, you could have for example 5 t and 5 kg of pure Ca with 4.69 x 1022 atoms of iodine as impurity. but if you are looking what amount of Ca there is if you have 4.69 x 1022 atoms of iodine in calcium iodine compound (CaI2) then it is another story! if you take that there is 4.69 x 1022 atoms of iodine in CaI2 compound you must notice that it states ATOMS not molecules of I2 but I atoms! then you notice that same number of Ca atoms is there, if you were asked for amount of I2 then you would have to divide given number with two cause there are 2 atoms for one molecule of I2, but to conclude you have: n=N/L= 4.69 x 1022/6,022*10^23 mol-1 = 0,0779 mole of Ca in sample!

OpenStudy (vincent-lyon.fr):

This question is nonsensical! "Pure calcium containing iodine" is contradictory.